OpenStudy (anonymous):

tan theta*sin theta=sec theta - cos theta, its proofs, but idk how to do them

OpenStudy (chaise):

well, tan(x)=sin(x)/cos(x) and sec(x)=1/cos(x). Does this help?

OpenStudy (anonymous):

so sin(x)/cos(x)*sin(x)=1/cos(x)*cos(x)?